a carpenter sells a centre table for rupees 322 , getting (1.6)th of its cost price . find his gain percent
Answer by stanbon(75887)

a carpenter sells a centre table for rupees 322 , getting (1.6)th of its cost price . find his gain percent
------
Let original cost be "x":
Equation:
1.6x = 322
x = 201.25
----
Quantity gain = 322-201.25 = 120.75
---
Percent gain = 120.75/201.25 = 0.6 = 60%
=============================================
Cheers,
Stan H.
==============